OpenStudy (anonymous):

find the equation of the line passing through the point (1,1/2) and perpendicular to the line y+4x=6

OpenStudy (anonymous):

\[y+4x=6\] \[y=-4x+6\] \[m=-4\] \[m_{p}=\frac{1}{4}\] \[y-y_{1}=m_{p}(x-x_{1})\] \[(x_{1}, y_{1})=(1,\frac{1}{2})\] \[y-\frac{1}{2}=\frac{1}{4}(x-1)\]

OpenStudy (anonymous):

m is the gradient of the line you were given. "mp" is the gradient perpendicular to that line.

OpenStudy (anonymous):

You can finish that off.

OpenStudy (anonymous):

@alberto123456 Mate, you going to finish that off or just keep bumping the thread?